### v4

@@ Abstract
We present a theoretical framework for the cuprate superconductors, rooted in a fractionalized Fermi liquid (FL*) description of the intermediate-temperature pseudogap phase at low doping. The FL* theory predicted hole pockets each of fractional area $p/8$ at hole doping $p$, in contrast to the area $p/4$ in a spin density wave state or its thermal fluctuation. A recent magnetotransport observation of the Yamaji angle is in good agreement with area $p/8$.
We review a theory for the FL* phase of a single-band model using a layer construction with a pair of ancilla qubits on each site: the Ancilla Layer Model (ALM). Its mean field theory yields hole pockets of area $p/8$, and matches the gapped photoemission spectrum in the anti-nodal region of the Brillouin zone. Fluctuations are described by the SU(2) gauge theory of a background spin liquid with critical Dirac spinons. A Monte Carlo study of the thermal SU(2) gauge theory transforms the hole pockets into Fermi arcs in photoemission. One route to confinement of FL* upon lowering temperature yields a $d$-wave superconductor via a Kosterlitz-Thouless transition of $h/(2e)$ vortices, with nodal Bogoliubov quasiparticles featuring anisotropic velocities and vortices surrounded by charge order halos. An alternative route produces a charge-ordered metallic state that exhibits quantum oscillations in agreement with experimental data.
Increasing doping from the FL* phase in the ALM drives a transition to a conventional Fermi liquid at large doping, passing through an intermediate strange metal regime. We formulate a theory of this metal using a critical quantum ‘charge’ liquid of mobile electrons in the presence of disorder, developed via an extension of the Sachdev-Ye-Kitaev model.
@@
This perspective connects the pseudogap, superconducting, charge-ordered, and strange metal phases within a single theoretical landscape.
@@ 1 Introduction
Figure 1: Cuprate phase diagram from Fig. 4 of Ref. [1]. (The only changes from the 2010 figure are that the pseudogap metal is now labeled FL*, dopings are denoted by $p$ rather than $x$, arrows have been removed, and section number annotations have been added; but all phase boundaries, crossovers, and critical points are the same.) The incipient FL*-FL transition at $p_{c}$ is that at $s_{c}$ in the normal state in Fig. 37; there is no transition at $p_{c}$ in the presence of superconductivity. The SDW transition at $p_{\rm sdw}$ is discussed in Section 9 in the presence of disorder, where the “spin gap” region above becomes a Griffiths phase.
@@
Already in 1987, Anderson [3] suggested that ‘resonating valence bond states’ hold the key to the physics of the cuprates. Today, there is a well-developed theory of such quantum spin liquid states with complex many-particle entanglement, which host a variety of well-defined particle-like anyonic excitations [4]. These spin liquids are analogs of fractional quantum Hall states, but often broken time-reversal symmetry. But, despite much effort, such states have not shed significant light on the cuprate phase diagram.
Instead, our focus here is on ‘critical quantum liquids’ which do not have any particle-like excitations. One such critical state is a particular quantum spin liquid, described in Section 3, in which the charge of some electrons is localized, while their spins remain entangled in a scale-invariant manner—it leads to a description of the pseudogap region of Fig. 2 and of its low temperature instabilities. Another distinct critical state is associated with the collective dynamics of electron charge, a critical quantum ‘charge’ liquid, characteristic of the strange metal regime of Figs. 1 and 2—its theoretical description builds on concepts from the Sachdev-Ye-Kitaev (SYK) model, introduced in Section 7.
@@ 1.1 Pseudogap
But this powerful paradigm runs into difficulty at lower $p$, and especially in the regime of the pseudogap, where there is no complete Fermi surface above $T_{c}$ upon which the Cooper pairs can form [8, 9, 10, 11, 12, 13, 14, 15, 16] (see Fig. 24a). Nevertheless, below $T_{c}$, there is no dramatic change in the nature of the superconductivity with varying $p$, and so it is clear that the superconductor itself is always adiabatically connected to a $d$-wave BCS state.
Recent magnetotransport measurements in the pseudogap regime [17, 18] reveal results that stand in striking contrast to earlier photoemission observations. Photoemission reveals a ‘Fermi arc’ in the nodal regime of the Brillouin zone, along with a dispering gapped spectrum in the anti-nodal region (see Fig. 12a). In contrast, magnetotransport is best modeled by complete hole pockets, and the observed Yamaji angle [18] is close to that obtained from hole pockets of area $p/8$. We shall present a theory in this paper which aims to resolve this tension between photoemission and magnetotransport.
@@
We describe the low $p$ phases by a ‘fractionalized Fermi liquid’ [36, 37, 38, 39, 40, 41] (FL*) theory of the pseudogap metal. The FL* is a state with Fermi surfaces which do not enclose the Luttinger area, and this is possible if there is a ‘background’ spin liquid, as is reviewed in Section 2.
@@
In Section 2.2 we discuss the more subtle theory needed to obtain a FL* phase as a mean-field solution in the single-band model. Such a mean-field solution can be achieved by introducing a layer construction with a pair of ancilla qubits [48] on each site, so that we have a three-layer Ancilla Layer Model (ALM); see Fig. 7, and the complete ancilla gauge theory for a single-band model in the Appendix. As we will describe, such a layer constructions allows exploration of a wider class of mean-field and gauge theories, and variational wavefunctions, in rough analogy to the utility of layered neural networks [49]. The mean-field theory of the single-band FL* is obtained by assuming that the top two layers form the familiar Kondo lattice heavy Fermi liquid FL phase—this has a spinon Fermi surface in the middle layer which hybridizes with a condensed Higgs (‘slave’) boson $\Phi$. Although the resulting Fermi surface has the Luttinger area for the Kondo lattice, it does not have the Luttinger area for a single-band Hubbard model. Reassuringly, in this mean-field, the bottom ancilla layer is a decoupled spin liquid which is required to exist by the non-Luttinger area. This leads to the slogan in Fig. 8 [50]
@@
Despite its simplicity, this mean-field theory can accurately model [51] the gapped electronic spectrum revealed by photoemission in the antinode in Fig. 12, as is discussed in Section 2.2.4. It also predicts pockets of area $p/8$ [36, 37, 52] consistent with magnetotransport, and distinct from the fluctuating spin density theory pocket area of $p/4$. The Yamaji angle observation of Chan et al. [18] is therefore direct evidence for the presence of fractionalization in the cuprates.
Furthermore, the mean-field theory leads to a variational wavefunction for FL* [48]: numerical studies [53, 54, 55] on such a wavefunction compare well with cold atom experiments on the square lattice fermionic Hubbard model [56], as we will discuss in Section 2.2.3.
But the main shortcoming of the mean-field theory of Section 2 is that the photoemission spectrum in the nodal region is not a Fermi arc.
@@
Sections 4 and 6 show that this critical spin liquid satisfies a key constraint: as the temperature is lowered, the emergent gauge fields are confined by a transition to a $d$-wave superconductor which is adiabatically connected to the BCS state [58]. The onset of superconductivity at low $p$ is not a BCS-type Cooper-pairing transition, but a confinement/Higgs transition of a gauge theory [47, 78]. This is a specific realization of an early suggestion by Anderson [3], that cuprate superconductivity appears by exploiting the pre-existing pairing of electrons in a resonating valence bond state.
There were several earlier proposals [79, 80, 81, 82, 57] for a transition from a spin liquid to a $d$-wave superconductor. In these theories, gapless fermionic spinons of the spin liquid directly transmute into the Bogoliubov quasiparticles of the $d$-wave superconductor with a massless Dirac dispersion. However, these proposals had a significant problem as they predict a nearly isotropic velocity dispersion, with the velocities along the Brillouin zone diagonals ($v_{F}$, see Fig. 26B) and the orthogonal direction ($v_{\Delta}$) being nearly equal to each other. Experimentally, we have $v_{F}/v_{\Delta}\sim 14$ to $19$ [83]. Section 6.1 shows how this problem is resolved [84, 85] by instead considering the transition from a FL* state to a $d$-wave superconductor. In such a theory, the spinons do not transmute into Bogoliubov quasiparticles; instead, they mutually annihilate with extraneous Bogoliubov quasiparticles from the ‘backside’ of the hole pockets.
@@
  • Section 5.1: The ‘Fermi arcs’ observed in photoemission [8, 9, 10, 14, 15, 16] and scanning tunneling microscopy (STM) [12, 13] are realized by thermal fluctuations of the SU(2) lattice gauge theory describing the FL* state [50]. The same computations also display quantum oscillations [50], which points to a reconciliation between photoemission and magnetotransport.
@@ 1.2 Strange metal
As indicated in Fig. 1 and 2, Sections 7 and 8 address the strange metal phase as a crossover between the pseudogap metal and the large $p$ Fermi liquid. Here a different critical quantum liquid plays an important role, one in which the electrons are mobile. Two solvable zero-dimensional models, the SYK model and the Yukawa-SYK model, yield much insight and their properties are reviewed in Section 7.
@@
We can imagine a situation, which we do not review here, where the continuous translational symmetry emerges only at very low temperatures, because of the dominance of umklapp processes. In this case, we can study zero-dimensional SYK-type solutions of the dynamic mean-field theory of a ‘Kondo breakdown’ quantum phase transition [92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106]. Such models have a non-vanishing entropy density in the zero temperature limit, and this has not been observed. However, we do expect that there is eventually a momentum conserving theory at low energies for clean lattice models [107], but it is possible that the crossover to such behavior is strongly suppressed in Kondo lattice models where the couplings between the Kondo spins are weak.
Instead, Section 8 examines the role of impurities, which are ubiquitous in experiments, and which will prevent the emergent continuous translational symmetry. Section 8.3 examines the generic situation where impurities lead to spatial randomness in the position of the quantum critical point. We describe a self-averaging treatment of the spatial disorder, inspired by the structure of the solution of the SYK model. This leads to a set of universal predictions at low temperatures ($T$) in the quantum-critical ‘fan’, which are largely independent of the particular quantum phase transition under consideration [90, 108]. These properties include a linear-in-$T$ resistivity, a $T\ln(1/T)$ specific heat, a $\sim 1/\omega$ tail in the optical conductivity at frequency $\omega$, and marginal Fermi liquid behavior in the electronic spectrum. These are in good agreement with observations across a wide range of correlated electron materials.

### v6

@@ Abstract
We review a theoretical framework for the cuprate superconductors, rooted in a fractionalized Fermi liquid (FL*) description of the intermediate-temperature pseudogap phase at low doping. The FL* theory predicted hole pockets each of fractional area $p/8$ at hole doping $p$, in contrast to the area $p/4$ in spin density wave theory. Magnetotransport measurements, including observation of the Yamaji angle, show clear evidence of hole pocket quasiparticles which can tunnel coherently between square lattice layers, and are consistent with the FL* description.
The FL* phase of a single-band model is described using a layer construction with a pair of ancilla qubits on each site: the Ancilla Layer Model (ALM). Its mean field theory yields hole pockets of area $p/8$, and matches the gapped photoemission spectrum in the anti-nodal region of the Brillouin zone. Fluctuations are described by the SU(2) gauge theory of a background spin liquid with critical Dirac spinons. A Monte Carlo study of the thermal SU(2) gauge theory transforms the hole pockets into Fermi arcs in photoemission. One route to confinement of FL* upon lowering temperature yields a $d$-wave superconductor via a Kosterlitz-Thouless transition of $h/(2e)$ vortices, with nodal Bogoliubov quasiparticles featuring anisotropic velocities and vortices surrounded by charge order halos. An alternative route yields a charge-ordered metallic state that has quantum oscillations consistent with observations.
Increasing doping from the FL* phase in the ALM drives a transition to a conventional Fermi liquid (FL) at large doping, passing through an intermediate strange metal regime. We formulate a theory of the FL*-FL metal-metal transition without a symmetry-breaking order parameter, using a critical quantum ‘charge’ liquid of mobile electrons in the presence of disorder, developed via an extension of the Sachdev-Ye-Kitaev model to two spatial dimensions.
@@ 1 Introduction
Figure 1: Cuprate phase diagram from Fig. 4 of Ref. [1]. (The only changes from the 2010 figure are that the pseudogap metal is now labeled FL*, dopings are denoted by $p$ rather than $x$, arrows have been removed, and section number annotations have been added; but all phase boundaries, crossovers, and critical points are the same.) The incipient FL*-FL transition at $p_{c}$ [2, 3, 4] is that at $s_{c}$ in the normal state in Fig. 37; there is no transition at $p_{c}$ in the presence of superconductivity. The SDW transition at $p_{\rm sdw}$ is discussed in Section 9 in the presence of disorder, where the “spin gap” region above becomes a Griffiths phase.
@@
Already in 1987, Anderson [6] suggested that ‘resonating valence bond states’ hold the key to the physics of the cuprates. Today, there is a well-developed theory of such gapped quantum spin liquid states with complex many-particle entanglement, which host a variety of well-defined particle-like anyonic excitations [7]. These gapped spin liquids are analogs of fractional quantum Hall states, but often without broken time-reversal symmetry.
There was much early effort on the structure of spin liquids doped away from the insulating half-filled limit by a density of $p$ holes [8]. Although there were significant advances, a number of difficulties have since become apparent. Two prominent difficulties are:
(i) In the underdoped ‘pseudogap’ regime (Fig. 2), angle-dependent magnetoresistance (ADMR) experiments [9, 10] provide evidence for small hole pockets which can tunnel coherently between the square lattice layers, and such coherent tunneling is not possible by fractionalized charge carriers in a doped spin liquid.
(ii) A $d$-wave superconductor with 4 nodal Bogoliubov quasiparticles is indeed obtained by doping a gapless spin liquid. But the observed velocities [11] of these quasiparticles are highly anisotropic, with the velocity perpendicular to the Fermi surface ($v_{F}$) being much larger than that in the orthogonal direction ($v_{\Delta}$).
In this review, we describe theoretical work on the fractionalized Fermi liquid (FL*) state [12, 13] as a model of the pseudogap metal, and as a starting point for a description of the rest of the cuprate phase diagram. We will describe how this approach addresses the difficulties noted above. As shown in Fig. 1, lowering temperature from FL* leads to a theory of the $d$-wave superconductor, and other observed states with charge order. Upon increasing doping from the pseudogap, we described the ‘strange metal’ by the quantum criticality of a quantum phase transition from FL* to a conventional Fermi liquid (FL) appropriate for the large doping regime. The FL*-FL transition [2, 3, 4] is an example of a metal-to-metal quantum phase transition without any associated symmetry breaking, unlike the well-known Hertz transitions [14] which require symmetry-breaking order parameters. Metal-to-metal transitions with increasing $p$ without any symmetry breaking have also appeared in numerical cluster dynamical mean-field studies by Tremblay and collaborators [15, 16, 17], and as an ‘orbital-selective Mott transition in momentum space’ [18, 19, 20, 21, 22, 23, 24]. We interpret the FL* state as the finite-dimensional description of the low doping metal of these works, properly accounting for the non-Luttinger Fermi surface.
The FL* state is characterized by an underlying quantum spin liquid, as we summarize in Section 1.1. For this spin liquid, we argue in Section 3 for a particular gapless, critical, quantum spin liquid, in which the charge of electrons is localized, while their spins are entangled in a scale-invariant manner. Unlike the gapped spin liquid states, this critical spin liquid has no particle-like excitations.
Another distinct critical state, without quasiparticle excitations, is associated with the FL*-FL transition and the collective dynamics of electron charge, as we summarize in Section 1.2. This is a critical quantum ‘charge’ liquid, characteristic of the strange metal regime of Figs. 1 and 2. Its theoretical description builds on concepts from the Sachdev-Ye-Kitaev (SYK) model, introduced in Section 7.
@@ 1.1 Pseudogap
But this powerful paradigm runs into difficulty at lower $p$, and especially in the regime of the pseudogap, where there is no complete Fermi surface above $T_{c}$ upon which the Cooper pairs can form, as shown in photoemission [28, 29, 30, 31, 32, 33, 34] and scanning tunneling microscopy (STM) [35, 36] experiments (see Fig. 24a). Nevertheless, below $T_{c}$, there is no dramatic change in the nature of the superconductivity with varying $p$, and so it is clear that the superconductor itself is always adiabatically connected to a $d$-wave BCS state.
Recent magnetotransport measurements in the pseudogap regime [9, 10] reveal results that stand in striking contrast to earlier photoemission observations. Photoemission reveals a ‘Fermi arc’ in the nodal regime of the Brillouin zone, along with a dispersing gapped spectrum in the anti-nodal region (see Fig. 12a). In contrast, magnetotransport is best modeled by complete hole pockets which can tunnel coherently between layers, and the observed Yamaji angle [10] is close to that obtained from hole pockets of area $p/8$. We shall review a theory in this paper which aims to resolve this tension between photoemission and magnetotransport.
@@
We describe the low $p$ phases by a ‘fractionalized Fermi liquid’ [54, 12, 55, 13, 56, 57, 2, 3, 58, 59, 60, 61, 62] (FL*) theory of the pseudogap metal. The FL* is a state with Fermi surfaces which do not enclose the Luttinger area, and this is possible if there is a ‘background’ spin liquid, as is reviewed in Section 2.
@@
While the existence of a FL* phase in a two-band Kondo lattice model is well-established, the existence of FL* in single-band Hubbard-type model was initially based on general theoretical arguments [69, 70, 2, 71, 72, 73], and a mapping to a quantum dimer model [74, 75]. But a systematic approach, based upon a mean-field theory plus fluctuations, appeared with the introduction of the Ancilla Layer Model [4]. As described in Section 2.2 and Appendix A, the ALM has a pair of ancilla qubits on each site, to yield a three-layer model: see Fig. 7. The complete ancilla gauge theory for a single-band model is described in Appendix B. Furthermore, the ALM can also describe the conventional FL phase, a feature which we shall exploit in Section 8.
The ALM allows exploration of a wider class of mean-field and gauge theories, and variational wavefunctions for the FL* state, in rough analogy to the utility of layered neural networks [76]. The mean-field theory of the single-band FL* is obtained by assuming that the top two layers form the familiar, and well established, Kondo lattice heavy Fermi liquid FL phase—this has a spinon Fermi surface in the middle layer which is unstable to hybridization with the conduction electrons with a condensed Higgs (‘slave’) boson $\Phi$, just as in the Kondo lattice model. But, although the resulting Fermi surface has the Luttinger area for the Kondo lattice, it does not have the Luttinger area for a single-band Hubbard model. Reassuringly, in this mean-field, the bottom ancilla layer is a decoupled spin liquid which is required to exist by the non-Luttinger area. This leads to the slogan in Fig. 8 (the nature of the spin liquid will be discussed in Section 3)
@@
Despite its simplicity, this mean-field theory can accurately model [77] the gapped electronic spectrum revealed by photoemission in the antinode in Fig. 12, as is discussed in Section 2.2.4. It also predicts pockets of area $p/8$ [12, 13, 69, 2, 78, 79] consistent with magnetotransport, and distinct from the fluctuating spin density theory pocket area of $p/4$. Moreover, the FL* state has coherent interlayer transport of electronic quasiparticles, while the spin density wave state requires inter-layer spin correlations, which are absent in HgBa2CuO4+δ [80, 81]. The Yamaji effect requires interlayer quasiparticle transport, and the Yamaji angle observations of Chan et al. [10] in HgBa2CuO4+δ are therefore direct evidence for the FL* state and the presence of fractionalization in the cuprates.
Furthermore, the mean-field theory leads to a variational wavefunction for FL* [4]: numerical studies [82, 83, 84] on such a wavefunction compare well with cold atom experiments on the square lattice fermionic Hubbard model [85, 86, 87], as we will discuss in Section 2.2.3.
But the main shortcoming of the mean-field theory of Section 2 is that the photoemission spectrum of the intermediate temperature pseudogap phase is not a Fermi arc in the nodal region.
@@
Sections 4 and 6 show that this critical spin liquid satisfies a key constraint: as the temperature is lowered, the emergent gauge fields are confined by a transition to a $d$-wave superconductor which is adiabatically connected to the BCS state [88]. The onset of superconductivity at low $p$ from FL* is not a BCS-type Cooper-pairing transition, but a confinement/Higgs transition of a gauge theory [88, 60, 108], as indicated in Fig. 1. This is a specific realization of an early suggestion by Anderson [6], that cuprate superconductivity appears by exploiting the pre-existing pairing of electrons in a resonating valence bond state. On the other hand, the larger $p$ transition to superconductivity from the FL is of the BCS-type. Note that the critical singularity of the thermal superconducting transition is always described by the Kosterlitz-Thouless theory of $h/(2e)$ vortices [108], and there is no singularity as a function of $p$ along the critical line. The distinction between small and large $p$ lies in the shorter-distance electronic structure, e.g. in the photoemission spectrum and the structure of the vortex core.
There were several earlier proposals [109, 110, 111, 112, 8] for a small $p$ transition from a spin liquid to a $d$-wave superconductor. In these theories, gapless fermionic spinons of the spin liquid directly transmute into the Bogoliubov quasiparticles of the $d$-wave superconductor with a massless Dirac dispersion. However, as noted earlier, these proposals had a significant problem as they predict a nearly isotropic velocity dispersion, with the velocities along the Brillouin zone diagonals ($v_{F}$, see Fig. 26B) and the orthogonal direction ($v_{\Delta}$) being nearly equal to each other. Experimentally, we have $v_{F}/v_{\Delta}\sim 14$ to $19$ [11]. Section 6.1 shows how this problem is resolved [113, 114] by instead considering the transition from a FL* state to a $d$-wave superconductor. In such a theory, the spinons do not transmute into Bogoliubov quasiparticles; instead, they mutually annihilate with extraneous Bogoliubov quasiparticles from the ‘backside’ of the hole pockets.
@@
  • Section 5.1: The ‘Fermi arcs’ observed in photoemission [28, 29, 30, 32, 33, 34] and STM [35, 36] are realized by thermal fluctuations of the SU(2) lattice gauge theory describing the FL* state [108]. The same computations also display quantum oscillations [108], which points to a reconciliation between photoemission and magnetotransport. We note that the conversion of the hole pockets into arcs is primarily a thermal effect. In the FL* state at $T=0$, a non-zero quasiparticle residue is present around the hole pocket even in the presence of quantum fluctuations of the SU(2) gauge theory [77].
@@ 1.2 Strange metal
As indicated in Fig. 1 and 2, Sections 7 and 8 address the strange metal phase as a crossover between the pseudogap metal and the large $p$ Fermi liquid, induced by an underlying metal-to-metal quantum phase transition without any symmetry breaking. Such a metal-metal transition appeared in the numerical cluster dynamic mean-field analysis of Sordi et al. [15, 16, 17] and is connected to the ‘orbital-selective Mott transition in momentum space’ [18, 19, 20, 21, 22, 23, 24]. We identify [2, 3] the transition with the FL*-FL transition.
Sections 7-9 will connect the physics of the strange metal to an underlying FL*-FL transition. Our discussion of strange metals will be brief, and the reader is refereed to other reviews [119, 120, 121]: our purpose here is to connect the strange metal regime of the cuprate phase diagram to that of the pseudogap in earlier sections.
A different critical quantum liquid plays an important role in the physics of the strange metal, one in which the electrons are mobile. Two solvable zero-dimensional models, the SYK model and the Yukawa-SYK model, yield much insight and their properties are reviewed in Section 7.
@@
Dynamic mean-field theories. We can imagine a situation, reviewed elsewhere [119] and briefly in Section 8.1, where the continuous translational symmetry emerges only at very low temperatures, because of the dominance of umklapp processes. In this case, we can study zero-dimensional SYK-type solutions of the dynamic mean-field theory of a ‘Kondo breakdown’ quantum phase transition [119, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137], or related cluster dynamic mean-field theories of single-band models [18, 19, 20, 21, 15, 16, 22, 23, 17, 24]. Such models can apply at higher temperature in the ‘bad metal’ regime [119]. Our focus, instead, is on behavior at low temperatures, when we expect that umklapp processes are suppressed, and the arguments reviewed in Section 8.2 preclude strange metal behavior at low temperature. (Similarly, Shastry’s extremely correlated Fermi liquid (ECFL) [138, 139] displays Fermi liquid transport at the lowest temperatures.) Furthermore, the dynamic mean-field models have a non-vanishing entropy density in the zero temperature limit, and this has not been observed. However, it is possible that the crossover to a momentum conserving theory [140] is strongly suppressed in Kondo lattice models where the couplings between the Kondo spins are weak.
Instead, Section 8 examines the role of impurities, which are ubiquitous in experiments, and which will prevent the emergent continuous translational symmetry. Section 8.3 examines the generic situation where impurities lead to spatial randomness in the position of the quantum critical point using a two-dimensional Yukawa-Sadhdev-Ye-Kitaev (2D-YSYK) model. We describe a self-averaging treatment of the spatial disorder, inspired by the structure of the solution of the SYK model. This leads to a set of universal predictions at low temperatures ($T$) in the quantum-critical ‘fan’, which are largely independent of the particular quantum phase transition under consideration [122, 141]. These properties include a linear-in-$T$ resistivity, a $T\ln(1/T)$ specific heat, a $\sim 1/\omega$ tail in the optical conductivity at frequency $\omega$, and marginal Fermi liquid behavior in the electronic spectrum. These are in good agreement with observations across a wide range of correlated electron materials. We also note a recent study of the Hall co-efficient of the 2D-YSYK model [142], which yields a superlinear Hall effect.